Shift report 20030513 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue May 13 2003 - 23:57:22 EDT

  • Next message: Apache: "Shift report 20030514 00:00-08:00"
    ------------------------------------------------------------
    Date: Tue May 13 23:57:22 2003, Shift: 16:00-24:00
    Supervisor: sjs, with: 
    Log book#14 pages 109 - 109
    
    Shift Summary:
    ------------------------------------------------------------
    We had about 1 1/2 hours of beam at the beginning of the shift, then nothing until a few minutes ago.
    
    It was discovered during the shift that the yellow beam profile was bad and probably has been bad for the past several days.  This has been fixed.  
    
    A series of radiation interlock failures kept us from getting beam for the rest of the shift.
    
    Noticed that the Si detectors were not showing a pedestal.  Looks like we lost one of the preamp low voltages, although this need to be checked with an IR access. Hiro has turned off the Si preamp low voltage supplies in the meantime.
    
    Data runs 8949 and 8950. Just started run 8951
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Tue May 13 2003 - 23:58:34 EDT